# Benchmarking Vision Foundation Models for Input Monitoring in Autonomous Driving

### v2

@@ D.2 Covariate Shift Experiments
APS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.66 0.62 0.80 0.66 0.63 0.84 0.65 0.61 0.81 0.78 0.74 0.62
RN101 0.68 0.65 0.80 0.69 0.68 0.79 0.76 0.73 0.65 0.80 0.75 0.57
RN50x4 0.70 0.66 0.78 0.71 0.68 0.78 0.74 0.70 0.71 0.81 0.77 0.59
RN50x16 0.66 0.62 0.79 0.67 0.63 0.82 0.69 0.65 0.73 0.73 0.67 0.69
RN50x64 0.66 0.62 0.78 0.68 0.64 0.79 0.71 0.66 0.76 0.77 0.73 0.67
ViT-B/32 0.75 0.72 0.64 0.77 0.75 0.62 0.79 0.75 0.58 0.86 0.82 0.49
ViT-B/16 0.73 0.69 0.71 0.74 0.71 0.78 0.78 0.75 0.57 0.86 0.83 0.48
ViT-L/14 0.80 0.76 0.58 0.83 0.80 0.60 0.84 0.82 0.49 0.86 0.82 0.46
DINO ViT-S/16 0.66 0.60 0.68 0.72 0.66 0.68 0.69 0.63 0.65 0.76 0.71 0.60
ViT-S/8 0.61 0.58 0.80 0.65 0.60 0.77 0.65 0.61 0.78 0.73 0.66 0.62
ViT-B/16 0.65 0.60 0.73 0.71 0.65 0.69 0.67 0.63 0.72 0.75 0.72 0.74
ViT-B/8 0.65 0.58 0.72 0.71 0.65 0.68 0.66 0.62 0.72 0.80 0.74 0.52
RN50 0.59 0.56 0.88 0.58 0.57 0.90 0.60 0.57 0.85 0.57 0.56 0.91
DINOv2 ViT-S/14 0.55 0.53 0.90 0.56 0.53 0.90 0.57 0.54 0.86 0.63 0.59 0.83
ViT-B/14 0.58 0.55 0.88 0.60 0.57 0.92 0.59 0.56 0.85 0.66 0.61 0.77
ViT-L/14 0.55 0.53 0.91 0.56 0.54 0.89 0.56 0.53 0.92 0.59 0.55 0.88
ViT-G/14 0.56 0.53 0.89 0.61 0.57 0.84 0.57 0.54 0.89 0.64 0.60 0.83
Grounding DINO SwinB 0.58 0.54 0.85 0.60 0.57 0.86 0.66 0.61 0.79 0.86 0.81 0.36
SwinT 0.57 0.54 0.87 0.59 0.56 0.86 0.69 0.62 0.72 0.84 0.77 0.42
@@
CLIP RN50 0.85 0.82 0.48 0.86 0.84 0.53 0.85 0.81 0.47 0.94 0.91 0.23
RN101 0.87 0.84 0.43 0.88 0.88 0.47 0.94 0.93 0.26 0.94 0.92 0.20
RN50x4 0.90 0.87 0.36 0.91 0.89 0.37 0.94 0.93 0.26 0.96 0.95 0.18
RN50x16 0.88 0.86 0.40 0.89 0.87 0.42 0.90 0.88 0.31 0.92 0.88 0.29
RN50x64 0.90 0.86 0.30 0.91 0.89 0.32 0.92 0.91 0.26 0.94 0.92 0.21
ViT-B/32 0.92 0.91 0.30 0.93 0.93 0.27 0.96 0.95 0.16 0.98 0.98 0.08
ViT-B/16 0.92 0.91 0.29 0.93 0.92 0.29 0.96 0.96 0.18 0.98 0.98 0.10
ViT-L/14 0.96 0.95 0.14 0.97 0.97 0.11 0.98 0.97 0.07 0.97 0.96 0.10
DINO ViT-S/16 0.87 0.81 0.31 0.93 0.90 0.20 0.92 0.89 0.24 0.95 0.94 0.19
ViT-S/8 0.77 0.70 0.50 0.83 0.78 0.45 0.83 0.78 0.44 0.89 0.84 0.30
ViT-B/16 0.85 0.78 0.34 0.91 0.88 0.22 0.89 0.85 0.31 0.89 0.88 0.34
ViT-B/8 0.84 0.77 0.34 0.91 0.87 0.24 0.88 0.85 0.35 0.97 0.96 0.09
RN50 0.73 0.68 0.70 0.70 0.67 0.79 0.73 0.67 0.73 0.65 0.63 0.84
DINOv2 ViT-S/14 0.66 0.60 0.72 0.69 0.62 0.73 0.70 0.63 0.66 0.82 0.76 0.52
ViT-B/14 0.69 0.63 0.73 0.71 0.66 0.77 0.74 0.68 0.66 0.84 0.77 0.48
ViT-L/14 0.62 0.57 0.85 0.62 0.58 0.86 0.63 0.59 0.84 0.69 0.62 0.70
ViT-G/14 0.64 0.58 0.76 0.71 0.64 0.72 0.68 0.62 0.73 0.80 0.74 0.54
Grounding DINO SwinB 0.84 0.79 0.38 0.89 0.86 0.32 0.95 0.94 0.17 1.00 1.00 0.00
SwinT 0.87 0.84 0.34 0.90 0.89 0.30 0.96 0.94 0.10 1.00 1.00 0.02
@@
CLIP RN50 0.80 0.81 0.67 0.79 0.80 0.77 0.85 0.86 0.54 0.89 0.90 0.42
RN101 0.84 0.85 0.62 0.84 0.86 0.69 0.89 0.91 0.57 0.92 0.93 0.43
RN50x4 0.78 0.79 0.71 0.77 0.79 0.80 0.85 0.86 0.58 0.88 0.87 0.51
RN50x16 0.80 0.81 0.70 0.79 0.80 0.73 0.83 0.84 0.62 0.87 0.88 0.55
RN50x64 0.79 0.79 0.65 0.81 0.81 0.63 0.82 0.83 0.62 0.85 0.82 0.47
ViT-B/32 0.84 0.84 0.60 0.83 0.84 0.62 0.91 0.91 0.47 0.93 0.94 0.37
ViT-B/16 0.84 0.85 0.52 0.83 0.84 0.60 0.89 0.90 0.43 0.94 0.95 0.35
ViT-L/14 0.87 0.86 0.49 0.88 0.89 0.48 0.90 0.91 0.45 0.94 0.93 0.31
DINO ViT-S/16 0.86 0.85 0.47 0.90 0.90 0.43 0.91 0.91 0.42 0.95 0.96 0.28
ViT-S/8 0.82 0.81 0.55 0.86 0.85 0.51 0.87 0.88 0.54 0.93 0.93 0.35
ViT-B/16 0.87 0.87 0.41 0.91 0.91 0.33 0.91 0.92 0.37 0.87 0.88 0.57
ViT-B/8 0.87 0.87 0.45 0.91 0.92 0.35 0.90 0.91 0.51 0.95 0.96 0.33
RN50 0.64 0.60 0.84 0.54 0.53 0.90 0.69 0.67 0.82 0.64 0.62 0.87
DINOv2 ViT-S/14 0.70 0.67 0.80 0.69 0.67 0.88 0.76 0.74 0.78 0.79 0.76 0.72
ViT-B/14 0.64 0.62 0.86 0.64 0.62 0.87 0.70 0.68 0.85 0.76 0.75 0.78
ViT-L/14 0.60 0.58 0.88 0.60 0.58 0.89 0.61 0.61 0.88 0.67 0.65 0.80
ViT-G/14 0.60 0.57 0.83 0.64 0.61 0.83 0.63 0.62 0.88 0.72 0.72 0.80
Grounding DINO SwinB 0.86 0.85 0.51 0.86 0.86 0.49 0.95 0.95 0.26 0.99 0.99 0.05
SwinT 0.85 0.84 0.49 0.84 0.84 0.55 0.94 0.94 0.25 0.99 0.99 0.07
Table 10: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Foggy Cityscapes [57] with different intensity values and Bravo-Synflare [49] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, OC-SVM, GMM, and NF methods. Best values in bold.
Tabs. 10, 11, 12 and 13 present comprehensive covariate shift experiments that extend the analysis introduced in the main paper. Specifically, Tabs. 10 and 11 demonstrate experimental results where the monitoring set $\mathcal{X}_{\mathrm{monitor}}$ is derived from the Cityscapes training set, while Tabs. 12 and 13 utilize the ACDC reference training set as $\mathcal{X}_{\mathrm{monitor}}$.
@@
APS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.76 0.72 0.59 0.75 0.71 0.64 0.75 0.71 0.64 0.85 0.82 0.43
RN101 0.83 0.81 0.50 0.84 0.82 0.51 0.84 0.82 0.51 0.92 0.91 0.25
RN50x4 0.85 0.82 0.48 0.85 0.82 0.48 0.85 0.82 0.48 0.93 0.92 0.24
RN50x16 0.94 0.93 0.25 0.94 0.94 0.22 0.94 0.94 0.22 0.96 0.95 0.17
RN50x64 0.99 0.99 0.03 1.00 1.00 0.01 1.00 1.00 0.01 1.00 1.00 0.01
ViT-B/32 0.82 0.79 0.53 0.82 0.80 0.50 0.82 0.80 0.50 0.91 0.89 0.31
ViT-B/16 0.90 0.87 0.34 0.91 0.89 0.31 0.91 0.89 0.31 0.94 0.93 0.20
ViT-L/14 0.89 0.85 0.31 0.92 0.90 0.27 0.92 0.90 0.27 0.94 0.91 0.22
DINO ViT-S/16 0.87 0.83 0.33 0.93 0.91 0.20 0.93 0.91 0.20 0.94 0.93 0.21
ViT-S/8 0.89 0.84 0.30 0.95 0.93 0.16 0.95 0.93 0.16 0.96 0.95 0.13
ViT-B/16 0.90 0.87 0.25 0.96 0.95 0.11 0.96 0.95 0.11 0.95 0.94 0.17
ViT-B/8 0.95 0.94 0.15 0.99 0.99 0.06 0.99 0.99 0.06 0.98 0.98 0.10
RN50 0.49 0.47 0.92 0.40 0.42 0.97 0.40 0.42 0.97 0.56 0.52 0.87
DINOv2 ViT-S/14 0.89 0.83 0.33 0.89 0.84 0.35 0.89 0.84 0.35 0.96 0.92 0.16
ViT-B/14 0.80 0.75 0.58 0.81 0.76 0.61 0.81 0.76 0.61 0.88 0.83 0.45
ViT-L/14 0.73 0.67 0.64 0.74 0.68 0.67 0.74 0.68 0.67 0.80 0.75 0.56
ViT-G/14 0.71 0.65 0.66 0.77 0.71 0.57 0.77 0.71 0.57 0.79 0.74 0.61
Grounding DINO SwinB 1.00 1.00 0.02 1.00 0.99 0.02 1.00 0.99 0.02 1.00 1.00 0.00
SwinT 0.99 0.99 0.02 0.99 0.99 0.03 0.99 0.99 0.03 1.00 1.00 0.00
Table 11: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Bravo-Synrain [49] dataset. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, OC-SVM, GMM, and NF methods. Best values in bold.
@@
APS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.95 0.94 0.22 0.94 0.93 0.22 0.98 0.98 0.07 0.97 0.94 0.09
RN101 0.94 0.94 0.25 0.94 0.93 0.28 0.98 0.97 0.12 0.98 0.96 0.09
RN50x4 0.97 0.97 0.13 0.97 0.96 0.16 0.99 0.98 0.07 0.98 0.98 0.06
RN50x16 0.97 0.97 0.11 0.97 0.97 0.12 0.98 0.98 0.08 0.98 0.96 0.08
RN50x64 0.98 0.98 0.07 0.99 0.99 0.06 0.98 0.98 0.06 0.99 0.98 0.04
ViT-B/32 0.97 0.97 0.12 0.97 0.97 0.11 0.99 0.99 0.03 0.99 0.99 0.04
ViT-B/16 0.95 0.95 0.20 0.95 0.95 0.18 0.98 0.98 0.07 0.98 0.98 0.07
ViT-L/14 0.98 0.97 0.09 0.98 0.97 0.09 1.00 0.99 0.02 0.99 0.98 0.03
DINO ViT-S/16 0.80 0.77 0.57 0.79 0.76 0.60 0.90 0.88 0.34 0.87 0.84 0.44
ViT-S/8 0.68 0.63 0.71 0.70 0.65 0.70 0.77 0.70 0.56 0.81 0.77 0.52
ViT-B/16 0.74 0.70 0.66 0.74 0.70 0.65 0.86 0.82 0.43 0.85 0.83 0.50
ViT-B/8 0.81 0.77 0.48 0.83 0.80 0.46 0.90 0.88 0.35 0.91 0.90 0.35
RN50 0.92 0.91 0.42 0.92 0.92 0.38 0.90 0.90 0.38 0.64 0.65 0.89
DINOv2 ViT-S/14 0.82 0.79 0.54 0.79 0.75 0.58 0.91 0.87 0.23 0.90 0.83 0.24
ViT-B/14 0.74 0.71 0.67 0.73 0.68 0.69 0.84 0.78 0.46 0.85 0.79 0.44
ViT-L/14 0.61 0.58 0.84 0.60 0.56 0.79 0.72 0.66 0.69 0.75 0.67 0.59
ViT-G/14 0.62 0.60 0.82 0.67 0.63 0.71 0.77 0.73 0.63 0.83 0.77 0.46
Grounding DINO SwinB 0.76 0.73 0.57 0.82 0.81 0.50 0.96 0.95 0.16 0.98 0.97 0.08
SwinT 0.80 0.79 0.54 0.83 0.83 0.52 0.98 0.97 0.08 0.96 0.95 0.13

### v3

@@ D.2 Covariate Shift Experiments
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.66 0.62 0.80 0.66 0.62 0.80 0.66 0.63 0.84 0.65 0.61 0.81 0.78 0.74 0.62
RN101 0.68 0.65 0.80 0.68 0.65 0.80 0.69 0.68 0.79 0.76 0.73 0.65 0.80 0.75 0.57
RN50x4 0.70 0.66 0.78 0.70 0.66 0.78 0.71 0.68 0.78 0.74 0.70 0.71 0.81 0.77 0.59
RN50x16 0.66 0.62 0.79 0.66 0.62 0.79 0.67 0.63 0.82 0.69 0.65 0.73 0.73 0.67 0.69
RN50x64 0.66 0.62 0.78 0.66 0.62 0.77 0.68 0.64 0.79 0.71 0.66 0.76 0.77 0.73 0.67
ViT-B/32 0.75 0.72 0.64 0.75 0.71 0.65 0.77 0.75 0.62 0.79 0.75 0.58 0.86 0.82 0.49
ViT-B/16 0.73 0.69 0.71 0.73 0.69 0.70 0.74 0.71 0.78 0.78 0.75 0.57 0.86 0.83 0.48
ViT-L/14 0.80 0.76 0.58 0.80 0.76 0.58 0.83 0.80 0.60 0.84 0.82 0.49 0.86 0.82 0.46
DINO ViT-S/16 0.66 0.60 0.68 0.66 0.60 0.68 0.72 0.66 0.68 0.69 0.63 0.65 0.76 0.71 0.60
ViT-S/8 0.61 0.58 0.80 0.61 0.58 0.80 0.65 0.60 0.77 0.65 0.61 0.78 0.73 0.66 0.62
ViT-B/16 0.65 0.60 0.73 0.65 0.60 0.73 0.71 0.65 0.69 0.67 0.63 0.72 0.75 0.72 0.74
ViT-B/8 0.65 0.58 0.72 0.65 0.58 0.72 0.71 0.65 0.68 0.66 0.62 0.72 0.80 0.74 0.52
RN50 0.59 0.56 0.88 0.59 0.56 0.88 0.58 0.57 0.90 0.60 0.57 0.85 0.57 0.56 0.91
DINOv2 ViT-S/14 0.55 0.53 0.90 0.55 0.53 0.90 0.56 0.53 0.90 0.57 0.54 0.86 0.63 0.59 0.83
ViT-B/14 0.58 0.55 0.88 0.58 0.55 0.88 0.60 0.57 0.92 0.59 0.56 0.85 0.66 0.61 0.77
ViT-L/14 0.55 0.53 0.91 0.55 0.53 0.91 0.56 0.54 0.89 0.56 0.53 0.92 0.59 0.55 0.88
ViT-G/14 0.56 0.53 0.89 0.56 0.53 0.89 0.61 0.57 0.84 0.57 0.54 0.89 0.64 0.60 0.83
Grounding DINO SwinB 0.58 0.54 0.85 0.57 0.54 0.84 0.60 0.57 0.86 0.66 0.61 0.79 0.86 0.81 0.36
SwinT 0.57 0.54 0.87 0.57 0.54 0.86 0.59 0.56 0.86 0.69 0.62 0.72 0.84 0.77 0.42
@@
CLIP RN50 0.85 0.82 0.48 0.85 0.82 0.47 0.86 0.84 0.53 0.85 0.81 0.47 0.94 0.91 0.23
RN101 0.87 0.84 0.43 0.87 0.84 0.44 0.88 0.88 0.47 0.94 0.93 0.26 0.94 0.92 0.20
RN50x4 0.90 0.87 0.36 0.90 0.87 0.36 0.91 0.89 0.37 0.94 0.93 0.26 0.96 0.95 0.18
RN50x16 0.88 0.86 0.40 0.88 0.86 0.40 0.89 0.87 0.42 0.90 0.88 0.31 0.92 0.88 0.29
RN50x64 0.90 0.86 0.30 0.90 0.86 0.31 0.91 0.89 0.32 0.92 0.91 0.26 0.94 0.92 0.21
ViT-B/32 0.92 0.91 0.30 0.92 0.90 0.30 0.93 0.93 0.27 0.96 0.95 0.16 0.98 0.98 0.08
ViT-B/16 0.92 0.91 0.29 0.92 0.91 0.29 0.93 0.92 0.29 0.96 0.96 0.18 0.98 0.98 0.10
ViT-L/14 0.96 0.95 0.14 0.96 0.95 0.15 0.97 0.97 0.11 0.98 0.97 0.07 0.97 0.96 0.10
DINO ViT-S/16 0.87 0.81 0.31 0.87 0.81 0.31 0.93 0.90 0.20 0.92 0.89 0.24 0.95 0.94 0.19
ViT-S/8 0.77 0.70 0.50 0.77 0.70 0.50 0.83 0.78 0.45 0.83 0.78 0.44 0.89 0.84 0.30
ViT-B/16 0.85 0.78 0.34 0.85 0.78 0.34 0.91 0.88 0.22 0.89 0.85 0.31 0.89 0.88 0.34
ViT-B/8 0.84 0.77 0.34 0.84 0.77 0.34 0.91 0.87 0.24 0.88 0.85 0.35 0.97 0.96 0.09
RN50 0.73 0.68 0.70 0.73 0.68 0.70 0.70 0.67 0.79 0.73 0.67 0.73 0.65 0.63 0.84
DINOv2 ViT-S/14 0.66 0.60 0.72 0.66 0.60 0.72 0.69 0.62 0.73 0.70 0.63 0.66 0.82 0.76 0.52
ViT-B/14 0.69 0.63 0.73 0.69 0.63 0.73 0.71 0.66 0.77 0.74 0.68 0.66 0.84 0.77 0.48
ViT-L/14 0.62 0.57 0.85 0.62 0.57 0.85 0.62 0.58 0.86 0.63 0.59 0.84 0.69 0.62 0.70
ViT-G/14 0.64 0.58 0.76 0.64 0.58 0.76 0.71 0.64 0.72 0.68 0.62 0.73 0.80 0.74 0.54
Grounding DINO SwinB 0.84 0.79 0.38 0.84 0.79 0.38 0.89 0.86 0.32 0.95 0.94 0.17 1.00 1.00 0.00
SwinT 0.87 0.84 0.34 0.87 0.84 0.34 0.90 0.89 0.30 0.96 0.94 0.10 1.00 1.00 0.02
@@
CLIP RN50 0.80 0.81 0.67 0.80 0.81 0.68 0.79 0.80 0.77 0.85 0.86 0.54 0.89 0.90 0.42
RN101 0.84 0.85 0.62 0.84 0.85 0.64 0.84 0.86 0.69 0.89 0.91 0.57 0.92 0.93 0.43
RN50x4 0.78 0.79 0.71 0.78 0.79 0.72 0.77 0.79 0.80 0.85 0.86 0.58 0.88 0.87 0.51
RN50x16 0.80 0.81 0.70 0.80 0.81 0.70 0.79 0.80 0.73 0.83 0.84 0.62 0.87 0.88 0.55
RN50x64 0.79 0.79 0.65 0.79 0.79 0.63 0.81 0.81 0.63 0.82 0.83 0.62 0.85 0.82 0.47
ViT-B/32 0.84 0.84 0.60 0.84 0.84 0.61 0.83 0.84 0.62 0.91 0.91 0.47 0.93 0.94 0.37
ViT-B/16 0.84 0.85 0.52 0.84 0.84 0.53 0.83 0.84 0.60 0.89 0.90 0.43 0.94 0.95 0.35
ViT-L/14 0.87 0.86 0.49 0.87 0.86 0.50 0.88 0.89 0.48 0.90 0.91 0.45 0.94 0.93 0.31
DINO ViT-S/16 0.86 0.85 0.47 0.86 0.85 0.47 0.90 0.90 0.43 0.91 0.91 0.42 0.95 0.96 0.28
ViT-S/8 0.82 0.81 0.55 0.82 0.81 0.55 0.86 0.85 0.51 0.87 0.88 0.54 0.93 0.93 0.35
ViT-B/16 0.87 0.87 0.41 0.87 0.87 0.41 0.91 0.91 0.33 0.91 0.92 0.37 0.87 0.88 0.57
ViT-B/8 0.87 0.87 0.45 0.87 0.87 0.45 0.91 0.92 0.35 0.90 0.91 0.51 0.95 0.96 0.33
RN50 0.64 0.60 0.84 0.64 0.60 0.84 0.54 0.53 0.90 0.69 0.67 0.82 0.64 0.62 0.87
DINOv2 ViT-S/14 0.70 0.67 0.80 0.70 0.67 0.80 0.69 0.67 0.88 0.76 0.74 0.78 0.79 0.76 0.72
ViT-B/14 0.64 0.62 0.86 0.64 0.62 0.86 0.64 0.62 0.87 0.70 0.68 0.85 0.76 0.75 0.78
ViT-L/14 0.60 0.58 0.88 0.60 0.58 0.88 0.60 0.58 0.89 0.61 0.61 0.88 0.67 0.65 0.80
ViT-G/14 0.60 0.57 0.83 0.60 0.57 0.83 0.64 0.61 0.83 0.63 0.62 0.88 0.72 0.72 0.80
Grounding DINO SwinB 0.86 0.85 0.51 0.86 0.85 0.51 0.86 0.86 0.49 0.95 0.95 0.26 0.99 0.99 0.05
SwinT 0.85 0.84 0.49 0.85 0.84 0.49 0.84 0.84 0.55 0.94 0.94 0.25 0.99 0.99 0.07
Table 11: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Foggy Cityscapes [59] with different intensity values and Bravo-Synflare [50] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
Tabs. 11, 12, 13 and 14 present comprehensive covariate shift experiments that extend the analysis introduced in the main paper. Specifically, Tabs. 11 and 12 demonstrate experimental results where the monitoring set $\mathcal{X}_{\mathrm{monitor}}$ is derived from the Cityscapes training set, while Tabs. 13 and 14 utilize the ACDC reference training set as $\mathcal{X}_{\mathrm{monitor}}$.
@@
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.76 0.72 0.59 0.76 0.72 0.60 0.75 0.71 0.64 0.85 0.82 0.43 0.95 0.92 0.16
RN101 0.83 0.81 0.50 0.83 0.81 0.50 0.84 0.82 0.51 0.92 0.91 0.25 0.94 0.93 0.17
RN50x4 0.85 0.82 0.48 0.85 0.82 0.48 0.85 0.82 0.48 0.93 0.92 0.24 0.95 0.93 0.14
RN50x16 0.94 0.93 0.25 0.94 0.93 0.25 0.94 0.94 0.22 0.96 0.95 0.17 0.97 0.96 0.10
RN50x64 0.99 0.99 0.03 0.99 0.99 0.03 1.00 1.00 0.01 1.00 1.00 0.01 0.99 0.99 0.04
ViT-B/32 0.82 0.79 0.53 0.82 0.79 0.54 0.82 0.80 0.50 0.91 0.89 0.31 0.96 0.95 0.13
ViT-B/16 0.90 0.87 0.34 0.90 0.87 0.34 0.91 0.89 0.31 0.94 0.93 0.20 0.97 0.95 0.12
ViT-L/14 0.89 0.85 0.31 0.89 0.85 0.32 0.92 0.90 0.27 0.94 0.91 0.22 0.97 0.94 0.09
DINO ViT-S/16 0.87 0.83 0.33 0.87 0.83 0.33 0.93 0.91 0.20 0.94 0.93 0.21 1.00 1.00 0.02
ViT-S/8 0.89 0.84 0.30 0.89 0.84 0.30 0.95 0.93 0.16 0.96 0.95 0.13 0.99 0.99 0.02
ViT-B/16 0.90 0.87 0.25 0.90 0.87 0.25 0.96 0.95 0.11 0.95 0.94 0.17 0.99 0.99 0.03
ViT-B/8 0.95 0.94 0.15 0.95 0.94 0.15 0.99 0.99 0.06 0.98 0.98 0.10 1.00 1.00 0.00
RN50 0.49 0.47 0.92 0.49 0.47 0.92 0.40 0.42 0.97 0.56 0.52 0.87 0.65 0.64 0.85
DINOv2 ViT-S/14 0.89 0.83 0.33 0.89 0.83 0.33 0.89 0.84 0.35 0.96 0.92 0.16 0.91 0.84 0.20
ViT-B/14 0.80 0.75 0.58 0.80 0.75 0.58 0.81 0.76 0.61 0.88 0.83 0.45 0.94 0.90 0.18
ViT-L/14 0.73 0.67 0.64 0.73 0.67 0.64 0.74 0.68 0.67 0.80 0.75 0.56 0.88 0.83 0.37
ViT-G/14 0.71 0.65 0.66 0.71 0.65 0.66 0.77 0.71 0.57 0.79 0.74 0.61 0.90 0.87 0.35
Grounding DINO SwinB 1.00 1.00 0.02 1.00 1.00 0.02 1.00 0.99 0.02 1.00 1.00 0.00 1.00 1.00 0.00
SwinT 0.99 0.99 0.02 0.99 0.99 0.02 0.99 0.99 0.03 1.00 1.00 0.00 1.00 1.00 0.00
Table 12: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Bravo-Synrain [50] dataset. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
@@
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.95 0.94 0.22 0.95 0.94 0.22 0.94 0.93 0.22 0.98 0.98 0.07 0.97 0.94 0.09
RN101 0.94 0.94 0.25 0.94 0.94 0.25 0.94 0.93 0.28 0.98 0.97 0.12 0.98 0.96 0.09
RN50x4 0.97 0.97 0.13 0.97 0.97 0.13 0.97 0.96 0.16 0.99 0.98 0.07 0.98 0.98 0.06
RN50x16 0.97 0.97 0.11 0.97 0.97 0.11 0.97 0.97 0.12 0.98 0.98 0.08 0.98 0.96 0.08
RN50x64 0.98 0.98 0.07 0.98 0.98 0.07 0.99 0.99 0.06 0.98 0.98 0.06 0.99 0.98 0.04
ViT-B/32 0.97 0.97 0.12 0.97 0.97 0.12 0.97 0.97 0.11 0.99 0.99 0.03 0.99 0.99 0.04
ViT-B/16 0.95 0.95 0.20 0.95 0.95 0.20 0.95 0.95 0.18 0.98 0.98 0.07 0.98 0.98 0.07
ViT-L/14 0.98 0.97 0.09 0.98 0.97 0.09 0.98 0.97 0.09 1.00 0.99 0.02 0.99 0.98 0.03
DINO ViT-S/16 0.80 0.77 0.57 0.80 0.77 0.57 0.79 0.76 0.60 0.90 0.88 0.34 0.87 0.84 0.44
ViT-S/8 0.68 0.63 0.71 0.68 0.63 0.71 0.70 0.65 0.70 0.77 0.70 0.56 0.81 0.77 0.52
ViT-B/16 0.74 0.70 0.66 0.74 0.70 0.66 0.74 0.70 0.65 0.86 0.82 0.43 0.85 0.83 0.50
ViT-B/8 0.81 0.77 0.48 0.81 0.77 0.48 0.83 0.80 0.46 0.90 0.88 0.35 0.91 0.90 0.35
RN50 0.92 0.91 0.42 0.92 0.91 0.42 0.92 0.92 0.38 0.90 0.90 0.38 0.64 0.65 0.89
DINOv2 ViT-S/14 0.82 0.79 0.54 0.82 0.79 0.54 0.79 0.75 0.58 0.91 0.87 0.23 0.90 0.83 0.24
ViT-B/14 0.74 0.71 0.67 0.74 0.71 0.67 0.73 0.68 0.69 0.84 0.78 0.46 0.85 0.79 0.44
ViT-L/14 0.61 0.58 0.84 0.61 0.58 0.84 0.60 0.56 0.79 0.72 0.66 0.69 0.75 0.67 0.59
ViT-G/14 0.62 0.60 0.82 0.62 0.60 0.82 0.67 0.63 0.71 0.77 0.73 0.63 0.83 0.77 0.46
Grounding DINO SwinB 0.76 0.73 0.57 0.75 0.72 0.58 0.82 0.81 0.50 0.96 0.95 0.16 0.98 0.97 0.08
SwinT 0.80 0.79 0.54 0.80 0.78 0.55 0.83 0.83 0.52 0.98 0.97 0.08 0.96 0.95 0.13